OpenStudy (zzr0ck3r):

Is 0 an even number? Is 1 a prime number? 0^0 = ?

OpenStudy (zzr0ck3r):

just some late night fun since its slow....

OpenStudy (anonymous):

- zero is dividable by 2 and either side is an odd number. - In the definition of a prime number the number must be >1 -0^0 is undefined and it is an indeterminate. ?

OpenStudy (anonymous):

Do you know of any methods other than differential calculus to answer mixing tank problems?

OpenStudy (vishweshshrimali5):

0^0 = 1 not undefined

OpenStudy (vishweshshrimali5):

Rest are correct.

OpenStudy (anonymous):

interesting could u explain please

OpenStudy (vishweshshrimali5):

Well it would be difficult and time consuming for me to type so I am just going to provide a reference link. One minute, let me search internet

OpenStudy (vishweshshrimali5):

Got one : https://www.math.hmc.edu/funfacts/ffiles/10005.3-5.shtml

OpenStudy (vishweshshrimali5):

Sorry, but feeling a bit lazy today ...

OpenStudy (anonymous):

lol, fair enough.

OpenStudy (zzr0ck3r):

0^0 is undefined

OpenStudy (vishweshshrimali5):

Lol I am stumped out.

OpenStudy (anonymous):

my calculator says it is but this webpage seems to prove otherwise

OpenStudy (zzr0ck3r):

0^0 = 1 implies ln(0^0) = ln(0) implies 0*ln(0) = 0 but ln(0 ) is undefined

OpenStudy (zzr0ck3r):

no that webpage says its undefined

OpenStudy (anonymous):

damn son, this is crazy

OpenStudy (vishweshshrimali5):

Not sure, then why these links says this: 1. http://mathforum.org/dr.math/faq/faq.0.to.0.power.html 2. http://www.askamathematician.com/2010/12/q-what-does-00-zero-raised-to-the-zeroth-power-equal-why-do-mathematicians-and-high-school-teachers-disagree/ But, correct @zzr0ck3r , read that again. Sorry

OpenStudy (anonymous):

this will make me sound dumb but how do i ask a question? I cant see the button.

OpenStudy (vishweshshrimali5):

*I meant I read that again. Sorry

OpenStudy (zzr0ck3r):

all those things are saying that it is undefined but it would be nice to call it 1 in some situations to make things easy. but just because it would be nice does not make it so. I already showed one example of the danger in doing such a thing

OpenStudy (vishweshshrimali5):

Yup...

OpenStudy (zzr0ck3r):

nono dont medal me for this lol

OpenStudy (vishweshshrimali5):

Medal for clearing doubt.

OpenStudy (vishweshshrimali5):

I always thought that it is safe to assume 0^0 = 1 for solving problems easily ;)

OpenStudy (zzr0ck3r):

NEVER AGAIN!

OpenStudy (zzr0ck3r):

jk

OpenStudy (vishweshshrimali5):

Lol

OpenStudy (vishweshshrimali5):

Not going to forget that again in my whole life :)

ganeshie8 (ganeshie8):

1/0 can be defined based on context, but not 0^0

OpenStudy (vishweshshrimali5):

limiting values ?

OpenStudy (zzr0ck3r):

yeah if we extend the real numbers

OpenStudy (zzr0ck3r):

I should have asked this in the meta-math section

OpenStudy (vishweshshrimali5):

Hmmm. Glad that you didn't, I never heard of that section before :D

OpenStudy (zzr0ck3r):

it might still post here actually. no idea.

OpenStudy (zzr0ck3r):

that could be a great section...

OpenStudy (vishweshshrimali5):

Hmm same here - No idea. Actually, I frequently discover some new sections here on OS